# VisBias: Measuring Explicit and Implicit Social Biases in Vision Language Models

### v2

@@ 5 Conclusions
This study has several limitations. (1) U.S.-centric demographic framing: We adopt the five racial labels Asian, Black, Hispanic, Middle Eastern, and White and rely on occupational imbalance statistics from the U.S. Bureau of Labor Statistics. These choices align with widely used U.S. taxonomies and facilitated data collection, yet they inevitably exclude many identities (e.g., Indigenous peoples, multiracial categories) and occupational patterns that prevail outside the U.S. (2) Selection of question options and attributes: In the explicit MCQ scenario we supply six income brackets, four education levels, four political orientations and four religions. By contrast, the implicit form completion task includes twenty attributes, eight education levels and six religions (including “None”). These design decisions were guided by (i) coverage in prior VLM bias work and (ii) the need to keep MCQs short enough to fit within model context limits. (3) The division of explicit and implicit biases: The distinction between MCQ/Yes-No tasks and image description/form completion reflects a pragmatic design choice. However, the evaluation of explicit and implicit biases is not confined to these tasks. For instance, MCQ can also be leveraged to revel implicit biases.
@@
This work investigates social biases in VLMs using publicly available images and model outputs. To elicit potential biases, we applied Caesar cipher-based prompts; these are not intended to generate harmful outputs but to surface otherwise hidden model behaviors. Human annotations are collected via Prolific with fair compensation and anonymized responses. Examples of biased outputs are included for transparency and research purposes only.
Appendix A Experiment Overview
@@
Appendix B Full Result Tables
@@ I E P R I E P R I E P R I E P R
Table 6: Stereotype score of descriptions generated by VLMs. It is the frequency (×$\times 1000$) of [GROUP]-related stereotype words in models’ descriptions about [GROUP] images. The highest scores across different VLMs are highlighted in bold, while those across genders/races are marked with underlines.
Appendix C Prompt for CipherChat
@@
Appendix D Available Choices for the 20 Attributes
@@
Appendix E Example Questionnaire
(a) Personality type.
(b) Religion.
@@ Appendix E Example Questionnaire
Appendix F Form Completion: Results for the Other 16 Attributes

### v3

@@ 6 Conclusions
This study has several limitations. (1) U.S.-centric demographic framing: We adopt the five racial labels Asian, Black, Hispanic, Middle Eastern, and White and rely on occupational imbalance statistics from the U.S. Bureau of Labor Statistics. These choices align with widely used U.S. taxonomies and facilitated data collection, yet they inevitably exclude many identities (e.g., Indigenous peoples, multiracial categories) and occupational patterns that prevail outside the U.S. (2) Selection of question options and attributes: In the explicit MCQ scenario we supply six income brackets, four education levels, four political orientations and four religions. By contrast, the implicit form completion task includes twenty attributes, eight education levels and six religions (including “None”). These design decisions were guided by (i) coverage in prior VLM bias work and (ii) the need to keep MCQs short enough to fit within model context limits. (3) The division of explicit and implicit biases: The distinction between MCQ/Yes-No tasks and image description/form completion reflects a pragmatic design choice. However, the evaluation of explicit and implicit biases is not confined to these tasks. For instance, MCQ can also be leveraged to revel implicit biases. (4) The diversity of image backgrounds: We purposely keep natural diversity (backgrounds, lighting, attire) to mirror in-the-wild usage where such cues are unavoidable. However, we acknowledge that uncontrolled visual artifacts may partly drive the observed associations. Future work can generate a Synthetic-VisBias subset with text-to-image models where every persona is rendered in an identical studio backdrop and neutral clothing palette.
@@
This work investigates social biases in VLMs using publicly available images and model outputs. To elicit potential biases, we apply Caesar cipher-based prompts; these are not intended to generate harmful outputs but to surface otherwise hidden model behaviors. Human annotations are collected via Prolific with fair compensation and anonymized responses. Examples of biased outputs are included for transparency and research purposes only.
LLM Usage
LLMs were employed in a limited capacity for writing optimization. Specifically, the authors provided their own draft text to the LLM, which in turn suggested improvements such as corrections of grammatical errors, clearer phrasing, and removal of non-academic expressions. LLMs were also used to inspire possible titles for the paper. While the system provided suggestions, the final title was decided and refined by the authors and is not directly taken from any single LLM output. In addition, LLMs were used as coding assistants during the implementation phase. They provided code completion and debugging suggestions, but all final implementations, experimental design, and validation were carried out and verified by the authors. Importantly, LLMs were NOT used for generating research ideas, designing experiments, or searching and reviewing related work. All conceptual contributions and experimental designs were fully conceived and executed by the authors.
Appendix A More Results
A.1 Experiment Overview
@@
A.2 Model Default Output without Images
Model MCQ w/o J MCQ w/ J Yes-No w/o J Yes-No w/ J
GPT-4o Refuse Income: All C
Education: All C
Political: All C
Religion: All B
All No All No
Gemini-1.5-Pro Refuse Refuse Refuse All No
LLaMA-3.2-Vision Refuse Refuse All No All No
Table 4: Model outputs when images are not given. “Refuse” denotes the response of “cannot provide answer without enough information.”
A.3 Quantitative Results
@@ I E P R I E P R I E P R I E P R
Table 8: Stereotype score of descriptions generated by VLMs. It is the frequency ($\times 1000$) of [GROUP]-related stereotype words in models’ descriptions about [GROUP] images. The highest scores across different VLMs are highlighted in bold, while those across genders/races are marked with underlines.
A.4 Form Completion: Results for the Other 16 Attributes
A.5 Form Completion: Results for Text Inputs
Model Independent Variable Value Dependent Variable KLD Human Verified Bias?
LLaMA Personality Type INTP Age 5.28 Yes
Veteran Status No Age 4.16 No
Physical Appearance Average height Race/Ethnicity 3.85 Yes
Gemini Language Spoken English Medical History 3.49 Yes
Occupation Student Age 3.14 No
Language Spoken English Occupation 2.98 Yes
GPT-4o Occupation Student Age 20.99 No
Education Level Doctorate Occupation 7.50 No
Residential Status Living with Family Occupation 6.90 No
Qwen- Hobbies/Interests Sport Occupation 4.97 Yes
Criminal Record No Age 3.92 Yes
Criminal Record No Disability Status 3.54 Yes
Table 9: Results (the largest three KL Divergence of each model) of (IV) Form Completion with text inputs. Models are: LLaMA-3.2-90B-Vision-Instruct Meta (2024), Gemini-1.5-Pro Team et al. (2024), GPT-4o Hurst et al. (2024), and Qwen-2.5-72B-InstructYang et al. (2024a).
Appendix B Detailed Settings
B.1 Prompt for CipherChat
@@
B.2 Randomly Sampled Examples of Collected Images
Figure 8: Examples of our collected images from Google Images.
B.3 Available Choices for the 20 Attributes
@@
Appendix C Human Study
C.1 More Details
Participants were compensated at an hourly rate of \$6.84, which is rated as “Good” by the Prolific platform, and each questionnaire required on average 5–10 minutes to complete. To ensure transparency and reproducibility, we further report detailed annotator demographics and task statistics. The average completion time was 455.46 seconds ($\approx$7.6 minutes, SD = 278.68). Among the 50 annotators, gender distribution was 24 female, 23 male, and 3 who preferred not to disclose. Reported race/ethnicity included 20 White, 16 Black, 7 Asian, and 7 Other/Not specified. The mean age was 31.32 years (SD = 8.00). In terms of employment status, 23 were employed full-time, 5 part-time, 8 unemployed, and 14 reported other or not specified. Annotators were primarily based in the United States (12), South Africa (12), Canada (7), the United Kingdom (3), and Italy (3). We also applied quality control filters to remove inattentive responses (e.g., under 2 minutes or uniform answers).
C.2 Example Questionnaire
@@
Figure 10: The example questions in our user study.